# MeSH Concept Relevance and Knowledge Evolution: A Data-driven Perspective

### v4

@@ 3.2 Top concepts in the MeSH terminology
Figure 3a displays the top 10 concepts according to the concept relevance, i.e., RRF, at level 2 (left side), level 3 (center), and level 4 (right). Half of the top concepts maintain their ranking over time. Concept G02 (Chemical Phenomena) shows variation among several positions, starting at the fourth position in 2014 and ending at the eighth in 2021. Meanwhile, C12 (Urogenital Diseases), D01 (Inorganic Chemicals), D02 (Organic Chemicals), and C01 (Infections) only change one position over time.
@@
In level 3, four concepts maintain a stable ranking, with B01.050 (Animals) leading. Concepts E05.481 (In Vitro Techniques), A03.620 (Liver), E01.789 (Prognosis), and N06.850 (Public Health) move among multiple positions, with ‘Prognosis’ tending to climb positions over time. In level 4, only the top three concepts, M01.060.116 (Adult), M01.060.406 (Child), and M01.060.057 (Adolescent), remain stable over time. These concepts continue to generate high research interest. The remaining concepts mainly decrease one position in the ranking.
Figure 4: Top level-3 concept relevance over time for categories (a) C, (b) D, and (c) E.
We computed the rank trend of concepts by adding the difference of two consecutive rank positions of a concept and dividing it by the number of differences, i.e., the number of years minus 1. Thus, we identified the concepts with the highest and lowest slopes. The top 10 upward concepts are shown in Figure 3b, which seem to lose positions over time at levels 2, 3, and 4. The behavior of the rank trend of concepts in these levels is similar. The concepts with the highest slope in level 2 are D06 (Hormones, Hormone Substitutes, and Hormone Antagonists), I01 (Social Sciences), D05 (Macromolecular Substances), and C24 (Occupational Diseases). The concept with the highest slope in level 3 is D01.930 (Thorium Compounds), and in level 4 is D03.383.517 (Oxathiins). The downwards concepts are shown in Figure 3c, which seems to gain positions over time, in contrast to the upward concepts. The concepts with the lowest rank trend slope are G16 (Biological Phenomena) and I02 (Education), both in level 2. In level 3, the concept M01.928 (Vegetarians) has the lowest trend slope, while in level 4, the concept E05.200.312 (COVID-19 Testing) has the lowest slope.
The bottom concepts are shown in Figure 3d. Concepts in level 2 remain stable, while in levels 3, and 4, concepts change positions constantly. On the other hand, bottom concepts in levels 3 and 4 change in a range of hundreds of positions, as in these levels, there are more concepts. Most of the changes occur towards the end of the studied period. The concepts J02.750 (‘Sustenance,’ in level 3) and B03.280.622 (‘Pannus,’ in level 4) seem to gain positions over time. In Figure 3d, right panel, we are only showing the plot since 2017 for visibility sake, as previous years follow the trend of 2017-2018. The MeSH headings of the concepts in Figure 3 can be found in Table 3, Table 4, and Table 5.
We analyzed the top concepts in concept relevance over time in categories C, D, and E at level 2. We mainly found no changes, except in category D, which presented a few changes. Thus, we show in Figure 4 the top concepts in these categories at level 3. In Figure 4a, four of the changing concepts belong to the branches of C12 (Urogenital Diseases) and C14 (Cardiovascular Diseases). The top concepts in Category D at Level 3 exhibit fewer changes than in Level 2. These concepts are D12.776 (Proteins), D12.644 (Peptides), and D12.125 (Amino Acids). In category E, E05.337 (Evaluation Studies as Topic) consistently loses positions over time, and E05.599 (Models, Theoretical) climbs some positions. Concepts E05.481 (In Vitro Techniques) and E01.789 (Prognosis) alternate in the top two positions. On the other hand, E05.581 (Methods) and E05.318 (Epidemiologic Methods) alternate the fourth and fifth positions. The MeSH headings of the concepts in Figure 4 can be found in Table 4.
@@ 3.3 Quantitative assessment
In this section, we describe the quantitative assessment of the multiple relevance aspects through the concept evolution task, and retracted papers. We used an ontology evolution dataset for the concept evolution task and identified retracted papers in PubMed. We start this section by describing the datasets, then detail each assessment and present their results.
@@ 3.3.1 Datasets for quantitative assessment
We created our dataset for concept evolution as previous studies still need to release their datasets. This dataset contains releases where MeSH descriptors are annotated with the type of change, ontological features, and data-driven temporal features. We calculated the latter features from the Data-properties dataset (explained in section 2.1).
A MeSH descriptor (or concept) has evolved if it has a change in the concept’s description (change of description), the addition of new children to the current concept (extension), a change of the concept’s location (move), or the removal of the concept (removal)  [14]. Table 2 shows the distribution of descriptors and changes over releases. We use the ontology versions of MeSH available in BioPortal, which have two versions per year. For example, in 2014, there were the 2014AA and 2014AB releases. When we downloaded the MeSH releases [^5] Downloaded on April 12, 2022, they were available from 2013 to 2021.
Table 2: Distribution of concepts over the MeSH releases from 2014 until 2021. The last two columns show the number of changed concepts in each release.
MeSH
@@
Evolved
Mean
relevance
2014AA 27,149 26,833 8.81E-04 316 2.53E-03
2014AB 27,455 27,341 8.73E-04 114 5.40E-03
2015AA 27,455 26,758 8.62E-04 697 2.04E-03
2015AB 27,883 27,879 8.80E-04 4 8.44E-04
2016AA 27,885 26,896 8.40E-04 989 1.96E-03
2016AB 28,472 28,472 8.65E-04 0 -
2017AA 28,472 27,673 8.35E-04 799 1.89E-03
2017AB 28,939 28,939 8.53E-04 0 -
2018AA 28,939 28,318 8.31E-04 621 1.88E-03
2018AB 29,352 29,349 8.43E-04 3 1.91E-03
2019AA 29,351 28,862 8.29E-04 489 1.67E-03
2019AB 29,640 29,640 8.36E-04 0 -
2020AA 29,640 29,279 8.20E-04 361 2.13E-03
2020AB 29,917 29,917 8.30E-04 0 -
2021AA 29,917 29,538 8.08E-04 379 2.51E-03
We used COntoDiff [37] to get the difference between two consecutive MeSH releases and found no changes in 2013. Thus, we excluded this year. We identified which concepts changed in each release and the change a concept has (change of description, extension, move, and removal). We used data until 2021AA, as we used 2021AA and 2021AB to identify the concepts that changed in 2021AA. Evolving concepts represent around 1% of concepts per release, and releases AB have no or very few changes.
Retracted papers in PubMed
We identified the retracted papers in the entire collection of PubMed 2022. Then, we identified which articles in the monthly citation networks (described in section 2.1) were retracted or non-retracted. On average, 800 papers were found to be retracted per month. We then retrieved the MeSH annotations of each article, along with the tree nodes associated with each MeSH descriptor. We obtained a value per metric for each article using these tree nodes by summing their metrics’ values.
Since our citation networks are based on a random sample, one paper may appear in consecutive months. Thus, we obtained an average of each metric per paper per year. We used these values to perform the statistical test.
3.3.2 Assessment in concept evolution
We focused on the concept evolution task to investigate whether the approach proposed in this work could contribute to biomedical terminology evolution. We performed a statistical test of significance using the Mann-Whitney test to determine if there is a significant difference between the relevance of descriptors that changed and descriptors that did not change among releases. We chose this non-parametric test as our data does not follow a normal distribution. For the concept relevance (RRF), we found that when there are evolving descriptors in a release, there is a statistical difference with a p-value less than 0.05, except for release 2015AB.
Figure 5a shows the test results for each relevance aspect in every release. Releases 2016AB, 2017AB, 2019AB, and 2020AB do not contain evolving descriptors. Consequently, more data is needed to compute the statistical test for these releases. Five out of eleven releases exhibit statistically significant differences in usefulness, while in influence and disruptiveness, eight show statistically significant differences. All releases show statistically significant differences in informativeness, but only one release in concept relevance does not demonstrate a statistical difference.
Release Disruptiveness Influence Informativeness Usefulness
Concept relevance
2014AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.03$ $p<0.001$
2014AB $p=0.25$ $p=0.13$ $p<0.001$ $p<0.001$ $p<0.001$
2015AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.64$ $p<0.001$
2015AB $p=0.53$ $p=0.56$ $p=0.03$ $p=0.35$ $p=0.29$
2016AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.005$ $p<0.001$
2016AB - - - - -
2017AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.11$ $p<0.001$
2017AB - - - - -
2018AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.22$ $p<0.001$
2018AB $p=0.07$ $p=0.07$ $p=0.04$ $p=0.06$ $p=0.04$
2019AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.31$ $p<0.001$
2019AB - - - - -
2020AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.001$ $p<0.001$
2020AB - - - - -
2021AA $p<0.001$ $p<0.001$ $p<0.001$ $p=0.002$ $p<0.001$
Figure 5: Evaluation of the multiple aspects in concept evolution. a) Results of the statistical significance test for each release per aspect using annotations of concept evolution. Green-coded cells correspond to a $p-value<0.05$; otherwise, it is red. b) The top evolving concepts according to each aspect are disruptiveness, influence, informativeness, and usefulness.
Figure 5b shows the concepts that mostly evolve according to each aspect. Some concepts change their position in disruptiveness and influence. Concepts of informativeness and usefulness remain mostly constant. The MeSH headings of concepts in Figure 5b can be found in Table 4.
@@ 3.3.3 Assessment in retracted papers
We used the multiple aspects metrics in this study to determine whether there is a significant difference between retracted and non-retracted papers in PubMed. We used the non-parametric Mann-Whitney test since our data does not follow a normal distribution. The results, presented in Figure 6a, revealed a significant difference between retracted and non-retracted papers in concept relevance, using the proposed aspects as the comparison unit, with a p-value less than 0.05. Further, when we performed the test per aspect, we found that there was also a statistical difference in usefulness. In influence, there is a significant difference in five out of eight years, while in disruptiveness and informativeness, seven show statistical differences.
Figure 6b shows the top concepts in retracted papers according to the rank relevance at level 3. Concepts B01.050 (Animals), A11.251 (Cells, Cultured), and E01.789 (Prognosis) lead the rank. Half of the concepts come from branch G, and three belong to branch G04 (Cell Physiological Phenomena). The MeSH headings of the concepts in Figure 6b can be found in Table 4.
@@
Figure 6: Evaluation of the multiple aspects in retracted papers. a) Results of the statistical significance test in retracted papers per aspect per year. Green-coded cells correspond to a $p-value<0.05$; otherwise, it is red. b) The top concepts in retracted papers according to the concept relevance at level 3.
@@ 4 Discussion
Our study focused on analyzing the evolution of MeSH using a data-driven approach. Leveraging concepts from information theory (i.e., entropy and category utility) and network analysis (i.e., disruptiveness and graph centrality), we investigated different aspects of relevance of MeSH concepts, such as informativeness, usefulness, disruptiveness, and influence. To summarize these four metrics in an overall concept relevance rank, we use the ranking fusion approach - RRF - from information retrieval. In our experiments, we analysed the relevance of MeSH concepts to the PubMed corpus, using the 2022 snapshot annotated with MeSH 2022. Using monthly data from 2014 to 2021, we compute the different relevance aspects, and show how they vary over time. We apply the methodology to two use-cases - concept evolution and retracted articles - and analysed the MeSH concept relevance in these scenarios.
@@
Overall, there is an increasing trend for informativeness and a decreasing trend for usefulness, influence, and disruptiveness. The latter is similar to a recent finding from a study of Park et al., which shows that disruptiveness in life and biomedicine sciences is decreasing  [36]. At the high level of the MeSH categories, our results showed categories C (Diseases)and D (Chemicals and Drugs) are the most relevant for the PubMed collection. While this result is somehow expected from the conceptual level, given the importance of such concepts in life and biomedical sciences, our method is able to quantify it according to different relevance aspects as well as an aggregated relevance rank. In particular, category D ranks first in terms of informativeness and usefulness while category C is the top category regarding influence and disruptiveness. When examining the metrics trends at level 2 of categories C, D, and E (Analytical, Diagnostic and Therapeutic Techniques, and Equipment), they behave similarly to level 1, showing mainly increasing trend informativeness and decreasing trends for the remaining aspects. The correlation analysis of descriptors revealed a weak correlation between informativeness, usefulness, and disruptiveness, demonstrating that indeed they capture different aspects of concept relevance.
@@
When analyzing the top 10 relevant concepts, higher levels tend to be more stable than lower levels due to the propagation of lower-level relevance scores. For example, level 1 appears more stable than level 2. Starting in level 2, i.e., for codes with 3 digits, we notice important changes in the relevance rank over time, with concepts gaining or losing several to many positions during the period analyzed. For example, the concept Chemical Phenomena (G02) falls 4 positions from 2014 to 2021. Levels 3 (6 digits) and 4 (9 digits) show an even more dynamic ranking. In level 3, Prognosis (E01.789) gains 3 positions, while Liver (A03.620) loses 3 positions in the same period. In level 4, most concepts in the top 10 lose one position over time. When examining the top concepts within categories C, D, and E, level 2 remains primarily stable. Among the top concepts within categories C and E at level 3, Cardiovascular Diseases (C14) and Investigative Techniques (E05) are the branches containing more changes, which might reflect changes around the scientific knowledge around those topics. Overall, we notice more concept relevance changes in lower ranks (bottom 10) than in upper ranks (top 10), apart from level 1.
@@
We further evaluated the application of our proposed approach using two analysis use-cases: concept evolution and retracted papers. For the former, we utilized an ontology evolution dataset, containing concept changes from consecutive MeSH versions. Interesting, the Mann-Whitney test confirmed a statistical difference between the relevance rank of MeSH concepts that evolve versus those that remain the same between version. This result suggest that the relevance of a concept could be used to identify candidates for concept update in KOS update. Additionally, our approach successfully discriminates between retracted and non-retracted papers even though the monthly proportion of retracted papers was approximately 0.0002%. In particular, concepts Animals (B01.050), Cells, Cultured (A11.251), Prognosis (E01.789), Age Groups (M01.060), Gene Expression Regulation (G05.308), and Biochemical Phenomena (G02.111) showed sustainable high-relevance to the retracted corpus. While the former 3 concepts also appear in the overall top 10 relevance list, the latter 3 concepts are particular of retracted papers, suggesting potential life and biomedical research areas with issues.
@@
This study has several limitations. First, as we use biomedical data sources, we must have a network of data instances using a KOS, i.e., PubMed and MeSH, to extend this work to other domains. Second, all the metrics were computed using a sample of the actual citation network. This was due to the algorithmic complexity of some of the metrics, which had a prohibitive computational time. Third, metrics are somehow correlated. Within a short period (one release to the other), our method is not sensitive enough to accurately identify all the changes between releases. Fourth, experiments in this work use PubMed 2022, which is annotated with MeSH 2022. More comprehensive experiments are needed to consider more MeSH versions.

### v5

@@ 3.2 Top concepts in the MeSH terminology
Note that some rank shifts reflect relative movements; small score differences can translate into large changes in rank due to the discretization of values. Figure 3a displays the top 10 concepts according to the concept relevance, at level 2 (left side), level 3 (center), and level 4 (right). Half of the top concepts maintain their ranking over time. Concept G02 (Chemical Phenomena) shows variation among several positions, starting in fourth place in 2014 and ending in eighth in 2021. Meanwhile, C12 (Urogenital Diseases), D01 (Inorganic Chemicals), D02 (Organic Chemicals), and C01 (Infections) change by one position over time.
@@
In level 3, four concepts maintain a stable ranking, with B01.050 (Animals) leading. Concepts E05.481 (In Vitro Techniques), A03.620 (Liver), E01.789 (Prognosis), and N06.850 (Public Health) move among multiple positions, with ‘Prognosis’ tending to climb positions over time. In level 4, only the top three concepts, M01.060.116 (Adult), M01.060.406 (Child), and M01.060.057 (Adolescent), remain stable over time. These concepts continue to generate high research interest. The remaining concepts mainly decrease by one position in the ranking.
We computed the rank trend of concepts by adding the differences between consecutive rank positions of a concept and dividing them by the number of differences, i.e., the number of years minus 1. Thus, we identified the concepts with the highest and lowest slopes. The top 10 upward concepts are shown in Figure 3b, which seem to gain positions over time at levels 2, 3, and 4. The behavior of the rank trend of concepts in these levels is similar. The concepts with the highest slope in level 2 are D06 (Hormones, Hormone Substitutes, and Hormone Antagonists), I01 (Social Sciences), D05 (Macromolecular Substances), and C24 (Occupational Diseases). The concept with the highest slope in level 3 is D01.930 (Thorium Compounds), and in level 4 is D03.383.517 (Oxathiins). The downward-trending concepts are shown in Figure 3c, which appear to lose positions over time, in contrast to the upward concepts. The concepts with the lowest rank trend slope are G16 (Biological Phenomena) and I02 (Education), both in level 2. In level 3, the concept M01.928 (Vegetarians) has the lowest trend slope, while in level 4, the concept E05.200.312 (COVID-19 Testing) has the lowest slope.
The bottom concepts are shown in Figure 3d. Concepts in level 2 remain stable, while in levels 3 and 4, concepts change positions constantly. On the other hand, bottom concepts in levels 3 and 4 vary more widely, often by hundreds of ranks, as in these levels, there are more concepts. Most of the changes occur towards the end of the studied period. The concepts J02.750 (‘Sustenance,’ in level 3) and B03.280.622 (‘Pannus,’ in level 4) seem to gain positions over time. In Figure 3d, right panel, we are only showing the plot since 2017 for clarity, as previous years follow the trend of 2017-2018. The MeSH headings of the concepts in Figure 3 can be found in the MeSH tree [34].
@@ 3.3 Quantitative assessment
In this section, we describe the quantitative assessment of the multiple relevance aspects through the concept evolution task and retracted papers. We used an ontology evolution dataset for the concept evolution task and identified retracted papers in PubMed. We first describe the datasets, then detail each assessment and present results. We chose the Mann-Whitney test as our data does not follow a normal distribution [29].
@@ 3.3.1 Datasets for quantitative assessment
We created our dataset for concept evolution as previous studies still need to release their datasets. This dataset contains releases where MeSH descriptors are annotated with the type of change and data-driven temporal features. We calculated these features from the Data-properties dataset (explained in section 2.1).
A MeSH descriptor (or concept) is considered to have evolved if it underwent a change in description (change of description), gained new children (extension), changed location (move), or was removed (removal)  [6]. We use the ontology versions of MeSH available in BioPortal, which have two versions per year, e.g., 2014AA and 2014AB releases. When we downloaded the MeSH releases [^6] Downloaded on April 12, 2022, releases were available from 2013 to 2021.
We used COntoDiff [20] to get the difference between two consecutive MeSH releases and found no changes in 2013. Thus, we excluded this year. We identified which concepts changed in each release and their type of change (description, extension, move, and removal). We used data until 2021AA, as we used 2021AA and 2021AB to identify the concepts that changed in 2021AA. Evolving concepts represent about 1% of concepts per release, and releases AB have no or very few changes. Over releases, there are $28,900$ concepts on average, the average relevance for evolving concepts is 2.09E032.09E-03, while for concepts that did not evolve is 8.46E048.46E-04.
Retracted papers in PubMed
We identified the retracted papers in the entire collection of PubMed 2022. Then, we identified the retracted papers in the monthly citation networks. On average, 800 papers were found to be retracted per month. We then retrieved the MeSH annotations of each article, along with the tree nodes associated with each MeSH descriptor. We obtained a value per metric for each article using these tree nodes by summing their metrics’ values. Since our citation networks are based on a random sample, a paper may appear in multiple months. We therefore averaged each metric per paper per year. We used these values to perform the statistical test.
3.3.2 Assessment in concept evolution
We focused on the concept evolution task to investigate whether the approach proposed in this work could contribute to biomedical terminology evolution. We performed a statistical test of significance using the Mann-Whitney test to determine if there is a significant difference between the relevance of descriptors that changed and descriptors that did not change among releases. For the concept relevance (RRF), we found statistical significant differences ($p<0.05$) whenever evolving descriptors were present.
Figure 4a shows the test results for each relevance aspect in every release. Releases 2016AB, 2017AB, 2019AB, and 2020AB do not contain evolving descriptors. Consequently, statistical tests could not be applied. Five out of eleven releases exhibit statistically significant differences in usefulness, while in influence and disruptiveness, eight show statistically significant differences. All releases show statistically significant differences in informativeness, but only one release in concept relevance does not demonstrate a statistical difference.
@@ 3.3.2 Assessment in concept evolution
2014AA $p<0.001$
2014AB $p<0.001$
2015AA $p<0.001$
2015AB $p=0.29$
2016AA $p<0.001$
2016AB
2017AA $p<0.001$
2017AB
2018AA $p<0.001$
2018AB $p=0.04$
2019AA $p<0.001$
2019AB
2020AA $p<0.001$
2020AB
2021AA $p<0.001$
Figure 4: Evaluation of the multiple aspects in concept evolution. a) Results of the statistical significance test for each release. Light gray-coded cells correspond to $p<0.05$; otherwise, it is dark gray. b) The top evolving concepts according to each aspect are (starting top left) disruptiveness, influence, informativeness, and usefulness.
Figure 4b shows the concepts that mostly evolve according to each aspect. Some concepts change their position in disruptiveness and influence. Concepts in informativeness and usefulness remain mostly constant. The concepts D12.776 (proteins), D12.644 (peptides), and D08.811 (enzymes) are among the top three aspects. The MeSH headings of the remaining concepts can be found in the MeSH tree [34].
@@ 3.3.3 Assessment in retracted papers
We used the multiple aspects metrics in this study to determine whether there is a significant difference between retracted and non-retracted papers in PubMed. The results (Figure 5a) show significant differences between retracted and non-retracted papers in concept relevance, using the proposed aspects as the comparison unit, with a $\textit{p}<0.05$. Furthermore, when we performed the test by aspect, we also found a statistical difference in usefulness. In influence, there is a significant difference in five out of eight years, while in disruptiveness and informativeness, seven years show differences.
Figure 5b shows the top concepts in retracted papers according to the rank relevance at level 3. Concepts B01.050 (Animals), A11.251 (Cells, Cultured), and E01.789 (Prognosis) lead the ranking. Half of the concepts come from branch G, and three belong to branch G04 (Cell Physiological Phenomena). The MeSH headings of the remaining concepts can be found in the MeSH tree [34].
@@
Figure 5: Evaluation of the aspects in retracted papers. a) Results of the statistical significance test. Light gray-coded cells correspond to $p<0.05$; otherwise, it is dark gray. b) Top-10 concepts in retracted papers according to the concept relevance at level 3.
@@ 4 Discussion
Previous research on MeSH has produced useful insights, though with limitations in capturing the multidimensional nature of concept relevance. Structural approaches of hierarchical growth patterns provide insight into organizational principles but do not reflect actual scientific impact or usage patterns [2]. In bibliometric and co-occurrence methods, researchers have attempted to bridge the gap between structure and usage  [7, 22, 24]. However, these approaches treat concepts as isolated units without considering citation relationships or hierarchical propagation.
Insights from the knowledge graph community offer complementary tools for ranking the importance of classes, i.e., concepts [12]. These approaches often overlook the rich citation dynamics available in bibliographic resources such as PubMed. Furthermore, studies on scientific disruptiveness indicate that citation patterns can reveal whether a topic displaces prior work or builds on it, insights that raw term counts fail to capture [43]. Finally, practical work on annotation and thesaurus maintenance highlights the need to prioritize which concepts to review or update in real-world applications [4].
A critical gap exists in instance-based analysis, where, despite millions of PubMed articles providing rich instantiation data for MeSH concepts, this resource has remained largely unexploited. Unlike previous work that relies on related or co-occurring descriptors, our approach leverages actual MeSH annotations as concept instantiations, providing a measurement of how concepts are employed in scientific discourse. By integrating information theory with network analysis and hierarchical propagation, our multidimensional framework addresses the limitations of single-perspective approaches, captures temporal evolution, and provides systematic validation through MeSH terminology changes and retracted articles.
@@
Leveraging concepts from information theory (i.e., entropy and category utility) and network analysis (i.e., disruptiveness and graph centrality), we investigated different aspects of the relevance of MeSH concepts. Data instances are propagated through direct mapping for information-theoretical measures and Algorithm 1 for network analysis metrics. Direct mapping considers instances when computing these measures. To summarize these four metrics in an overall concept relevance ranking, we applied the RRF method. In our experiments, we analysed the relevance of MeSH concepts to the PubMed corpus, using the 2022 snapshot annotated with MeSH 2022. Using monthly data from 2014 to 2021, we computed the relevance aspects and showed how they vary over time. We applied the methodology to two use-cases — concept evolution and retracted articles — and analyzed the MeSH concept relevance in these scenarios.
@@
Overall, there is an increasing trend for informativeness and a decreasing trend for usefulness, influence, and disruptiveness. The latter is similar to a recent finding from a study of Park et al., which shows that disruptiveness in life and biomedicine sciences is decreasing  [36]. At the high level of the MeSH categories, our results show that categories C (Diseases) and D (Chemicals and Drugs) are the most relevant for the PubMed collection. While this result is expected, given the importance of such concepts in life and biomedical sciences, our method can quantify their relevance according to different aspects, as well as an aggregated relevance rank. In particular, category D ranks first in terms of informativeness and usefulness, while category C leads in influence and disruptiveness. The correlation analysis of descriptors revealed a weak correlation among informativeness, usefulness, and disruptiveness, demonstrating that indeed they capture different aspects of concept relevance.
@@
Analysis of the top 10 relevant concepts shows that higher levels are more stable than lower ones due to the propagation of lower-level relevance scores; for instance, level 1 is more stable than level 2. At level 2 (three-digit codes), concepts undergo notable ranking shifts, with Chemical Phenomena (G02) falling by four positions between 2014 and 2021. Levels 3 (six digits) and 4 (nine digits) are even more dynamic: Prognosis (E01.789) rises three positions, while Liver (A03.620) drops three in the same period, and most level 4 concepts lose one position. Within categories C, D, and E, level 2 remains largely stable, but at level 3, Cardiovascular Diseases (C14) and Investigative Techniques (E05) show the greatest changes, possibly reflecting evolving scientific knowledge. Overall, lower-ranked concepts (bottom 10) shift more than higher-ranked ones, except at level 1.
@@
We further evaluated the application of our proposed approach using two analysis use-cases: concept evolution and retracted papers. For the former, we utilized an ontology evolution dataset, containing concept changes from consecutive MeSH versions. Interestingly, the Mann-Whitney test confirmed statistically significant differences between the relevance rank of MeSH concepts that evolve versus those that remain stable. This result suggests that the relevance of a concept could be used to identify candidates for update in KOS maintenance. Additionally, we show that retracted and non-retracted papers can be differentiated using our approach even though the monthly proportion of retracted papers was approximately 0.0002%. In particular, concepts Animals (B01.050), Cells, Cultured (A11.251), Prognosis (E01.789), Age Groups (M01.060), Gene Expression Regulation (G05.308), and Biochemical Phenomena (G02.111) showed consistently high relevance to the retracted corpus. While the former three concepts also appear in the overall top 10 relevance list, the latter three concepts are particular to retracted papers, suggesting potential life and biomedical research areas with issues.
@@
This study has several limitations. First, extending this work to other domains requires a dataset of instances annotated with a KOS, such as PubMed with MeSH. Alternatively, information extraction methods could be used to automatically annotate instances or corpus [8, 9, 31]. Second, all metrics were computed using a sample from the citation network due to the algorithmic complexity of some metrics, which made full computation prohibitive. Finally, metrics are somewhat correlated. Within a short period (e.g., one release to the next), our method is not always sensitive enough to identify all changes.